OpenStudy (anonymous):

I'm supposed to graph: y^2=x-9, so I changed it to y=radical x-9 and I graphed it on my graphing calculator by pressing the "y="'s button, putting in the equation and then pressing "graph", but I am not getting the correct graph. What's the problem?

OpenStudy (anonymous):

i answered you down there ..

OpenStudy (anonymous):

Did u not get my response?

OpenStudy (anonymous):

sorry now i read it.. what error do you get ?

OpenStudy (anonymous):

No error, but you need to plot 2 graphs.

OpenStudy (anonymous):

ERR: SYNTAX 1: QUIT 2: GOTO is what i get on my screen

OpenStudy (anonymous):

syntax error? what did you type? just plot these 2: \[y_1 = \sqrt{x-9}\] \[y_2 = -\sqrt{x-9}\]

OpenStudy (anonymous):

what do u get? my graph looks weird and is not one of my multiple choice options.

OpenStudy (anonymous):



OpenStudy (anonymous):

thats what i get

OpenStudy (anonymous):

You may need to adjust the window settings of the calculator.

OpenStudy (anonymous):

you should get :